# Factor Risk Budgeting and Beyond

### v1

@@ 5.1 Analysis of risk-based portfolios under an equity factor model
In this section, we construct various risk-based portfolios using the BARRA US Equity (USE3) model. [^11] A detailed description of the model can be found online or requested through https://www.msci.com/www/research-report/united-states-equity-model-use3/015920189. The USE3 is a multi-factor model covering 67 factors: 54 industry factors and 13 style factors (e.g. momentum, growth and size). The covariance matrix of the factor returns is provided, revealing the estimated relationship between the factors. The model covers a wide range of US equities, and provides the factor loading of each stock on each risk factor, as well as the idiosyncratic components for individual assets. [^12] It is worth noting that they implement a multi-industry approach, which means that stocks can have loadings on multiple industry factors. This approach enhances the model. It therefore corresponds to having a factor model of the following form.
$\Sigma=\beta\Sigma_{f}\beta^{\prime}+\Sigma_{idio}$
@@
We consider the stocks in the S&P500 index as of 31/10/2023 as our investment universe. [^13] The index actually has 503 components, as three of them have two share classes listed. We filter them and end up with a universe of 500 assets. We also use the parameters of the BARRA model for the same date. We calculate two Minimum Variance portfolios, one without constraint and another one with a long-only constraint, denoted by MV and $\text{MV}^{+}$ respectively. We also compute an RB portfolio (solving Problem (1)) where the risk budgets for each asset are equal – called the ERC portfolio. To introduce factor risk sensitivity into the portfolio construction process, we consider three different portfolios. The first is an FRB portfolio, in which factor risk budgets are chosen to be equal and computed by solving Problem (4.1). We call it the Equal Factor Risk Contribution portfolio, EFRC. The second portfolio is similar, but we restrict the weights to be positive following the ideas presented in Section 4.2. We denote it by $\text{EFRC}^{+}$. The third portfolio we consider, called the Equal Asset-Factor Risk Contribution portfolio, EAFRC, aims to find a balance between factor risk contributions and asset risk contributions as formalized by Problem (6) where we use equal asset and factor risk budgets and the following importance parameters: $\lambda_{a}=0.3$ and $\lambda_{f}=0.7$. Finally, we also include the equal-weighted portfolio, EW, in our analysis which is often a competitive challenger of risk-based methods.
Figure 1: Annualized ex-ante volatilities (left) and latent space representations (right) of the portfolios
@@
The results of the same analysis on factor risk contributions (using Proposition 4) are depicted in Figure 3. This time, EFRC plays the benchmark role as the least-concentrated portfolio in terms of factor risk contributions, distributing risk among factors equally. $\text{EFRC}^{+}$ and EAFRC follow this portfolio, providing factor risk distributions that are close to factor risk budgets. EW, ERC, $\text{MV}^{+}$ and MV are portfolios that do not allow factor risk contributions to be controlled by definition, and can result in portfolios where risk is concentrated on a small set of factors.
@@
In short, we see that the EW and ERC portfolios are well-balanced in terms of asset weights and asset risk contributions, while the Lorenz curves for factor risk contributions indicate a potential concentration of factor risk. The EFRC portfolio achieves perfectly an equal allocation of factor risk, but may suffer from a concentration of asset weights and asset risk contributions, as does the $\text{EFRC}^{+}$ portfolio. However, the EAFRC portfolio is, as expected, interesting for balancing asset and factor risk contributions. It seems to be the only strategy in our case that seeks a balance between all three attributes. Consequently, Asset-Factor Risk Budgeting portfolios may be strong candidates for managing the relationship between the three different pillars of portfolio risk management.
5.2 Constructing diversified multi-asset portfolios
In this section, we illustrate a case where we construct multi-asset portfolios within the framework of a macroeconomic factor model. Different asset classes can be affected by common macroeconomic factors, and it can therefore be interesting to manage the risk exposure to these factors when managing a multi-asset portfolio.
@@ 5.2 Constructing diversified multi-asset portfolios
Our objective is to investigate two strategies, namely Risk Budgeting and Asset-Factor Risk Budgeting portfolios where Expected Shortfall is chosen as the risk measure. More specifically, we want to analyze the evolution of asset and factor risk contributions over time by performing a backtest where we rebalance our portfolio on a weekly basis to obtain the desired portfolios based on the new information.
@@
First, we look at the case of the RB portfolio. The risk budgets are chosen as
@@
The choice of risk budgets depends on the portfolio manager. In our case, we assign 40% of risk budgets to bonds and equities, and 20% to commodities. We then allocate these budgets equally within each class. We find the RB portfolios for Expected Shortfall with a threshold $\alpha=95\%$ by solving Problem (2) using daily returns in the look-back window.
Figure 4: Evolution of asset and factor risk contributions (in %) for Risk Budgeting portfolios
In Figure 4, we observe that the risk contributions of the assets are very well aligned with the given vector of risk budgets, with the exception of small deviations which are linked to the accuracy of the Expected Shortfall estimates at a relatively high threshold $\alpha$ on a sample with a limited number of points. However, when we examine the factor risk contributions, which the RB portfolio does not take into account when optimizing portfolios, we observe that the risk contributions of Real Rates can reach levels of the order of 70%, which should preferably be avoided. On the contrary, the risk contributions from Broad Commodities are often marginal, despite the fact that 20% of risk budgets have been allocated to the commodity asset class through asset risk budgets.
To address this issue, we can consider Asset-Factor RB portfolios that allow us to control factor risk contributions as well as asset risk contributions. We choose the following factor risk budgets for Growth, Real Rates, Broad Commodities and Emerging Markets respectively:
@@
We consider the same asset risk budgets as in the case of the RB portfolio. We also need to set the asset and factor importance parameters, which we choose equal: $\lambda_{a}=0.5$ and $\lambda_{f}=0.5$. We can now compute the Asset-Factor RB portfolio by solving Problem (7) for Expected Shortfall with a threshold $\alpha=95\%$.
Figure 5: Evolution of asset and factor risk contributions (in %) for Asset-Factor Risk Budgeting portfolios
In Figure 5, we observe that the Asset-Factor RB portfolio maintains a balance between asset and factor risk contributions. Compared with the case of RB, we observe a deviation from asset risk budgets. However, this is the price we have to pay to stay close to factor risk budgets. We note that the risk contribution of Broad Commodities is no longer small, and that we achieve a significant exposure to this factor. Growth and Real Rates now have balanced contributions, in line with factor risk budgets. Thanks to the importance parameters, we can choose to be more conservative in one of the two types of risk contributions, depending on the investment objective or market conditions. In short, Asset-Factor RB portfolios can provide a flexible framework for better risk management of investment portfolios.
@@ 6 Conclusion
In this paper, we address the question of incorporating factor risk contributions into the portfolio optimization process. First, we introduce the concept of Risk Budgeting, which takes into account the risk contributions of individual assets. Next, we develop the necessary tools and propose a factor risk measure that allows us to apply these ideas to underlying risk factors in the presence of a linear factor model. We then formally define the Factor Risk Budgeting and an associated stochastic optimization problem for finding FRB portfolios for a large set of risk measures. We also propose a framework that combines RB and FRB to obtain portfolios that control the risk induced by both assets and factors. We conclude the paper with two case studies where we construct portfolios based on the mentioned ideas using equity and macroeconomic factor models.

### v2

@@ 6.1 Analysis of risk-based portfolios in an equity factor model
In this section, we construct various risk-based portfolios using the BARRA US Equity (USE3) model. [^16] A detailed description of the model can be found online or requested through https://www.msci.com/www/research-report/united-states-equity-model-use3/015920189. The USE3 is a multi-factor model covering 67 factors: 54 industry factors and 13 style factors (e.g. momentum, growth and size). The covariance matrix of the factor returns is provided, revealing the estimated relationship between the factors. The model covers a wide range of US equities, and provides the factor loading of each stock on each risk factor, as well as the idiosyncratic components for individual assets. [^17] It is worth noting that they implement a multi-industry approach, which means that stocks can have loadings on multiple industry factors. This approach enhances the model. It therefore corresponds to having a factor model of the following form:
$\Sigma=\beta\Sigma_{f}\beta^{\prime}+\Sigma_{idio},$
@@
We consider the stocks in the S&P500 index as of 31/10/2023 as our investment universe. [^18] The index actually has 503 components, as three of them have two share classes listed. We filter them and end up with a universe of 500 assets. We also use the parameters of the BARRA model for the same date. We calculate two Minimum Variance portfolios, one without constraint and another one with a long-only constraint, denoted by MV and $\text{MV}^{+}$ respectively. We also compute an RB portfolio (solving Problem (1)) where the risk budgets for each asset are equal – called the ERC portfolio. To introduce factor risk sensitivity into the portfolio construction process, we consider three different portfolios. The first is an FRB portfolio, in which factor risk budgets are chosen to be equal and computed by solving Problem (4.1). We call it the Equal Factor Risk Contribution portfolio, EFRC. The second portfolio is similar, but we restrict the weights to be positive following the ideas presented in Section 4.2. In other words, it is the portfolio obtained by solving Problem (4.1) under the constraint associated with the set $\mathcal{C}^{\geq 0}$. We denote it by $\text{EFRC}^{+}$. We also consider the Equal Asset-Factor Risk Contribution portfolio, EAFRC which aims to find a balance between factor risk contributions and asset risk contributions as formalized by Problem (6) where we use equal asset and factor risk budgets and the following importance parameters: $\lambda_{a}=0.3$ and $\lambda_{f}=0.7$. Finally, we also include the equally-weighted portfolio, EW, in our analysis which is often a competitive challenger of risk-based methods.
Figure 2: Annualized ex-ante volatilities (left) and latent space representations (right) of the portfolios.
@@
The results of the same analysis on factor risk contributions (using Proposition 4) are depicted in Figure 4. This time, EFRC plays the benchmark role as the least-concentrated portfolio in terms of factor risk contributions, distributing risk across factors equally. $\text{EFRC}^{+}$ and EAFRC follow this portfolio, providing factor risk distributions that are close to factor risk budgets. EW, ERC, $\text{MV}^{+}$ and MV are portfolios that do not allow factor risk contributions to be controlled by definition, and can result in portfolios where risk is concentrated on a small set of factors.
@@
In short, we see that the EW and ERC portfolios are well-balanced in terms of asset weights and asset risk contributions, while the Lorenz curves for factor risk contributions indicate a potential concentration of factor risk. The EFRC portfolio achieves perfectly an equal allocation of factor risk, but may suffer from a concentration of asset weights and asset risk contributions, as does the $\text{EFRC}^{+}$ portfolio. However, the EAFRC portfolio is, as expected, interesting for balancing asset and factor risk contributions. It seems to be the only strategy in our case that achieves a balance across all three attributes. Consequently, Asset-Factor Risk Budgeting portfolios may be strong candidates for managing the relationship between the three different pillars of portfolio risk management.
6.2 Constructing diversified cross-asset portfolios
In this section, we illustrate a case where we construct cross-asset portfolios within the framework of a macroeconomic factor model. Different asset classes can be affected by common macroeconomic factors, and it can therefore be interesting to manage the risk exposure to these factors when managing a cross-asset portfolio.
@@ 6.2 Constructing diversified cross-asset portfolios
Our objective is to investigate three strategies, namely Risk Budgeting, long-only Factor Risk Budgeting and Asset-Factor Risk Budgeting, where Expected Shortfall is chosen as the risk measure. More specifically, we want to analyse the performance of these strategies, but especially the evolution of asset and factor risk contributions over time by performing a backtest where we rebalance our portfolio on a weekly basis to obtain the desired portfolios based on the new information.
@@
First, we look at the case of the RB strategy. The risk budgets are chosen as
@@
The choice of risk budgets depends on the portfolio manager. In our case, we assign 40% of risk budgets to bonds and equities, and 20% to commodities. We then allocate these budgets equally within each class. The strategy corresponds to finding the RB portfolio for Expected Shortfall with a threshold $\alpha=95\%$ by solving Problem (2) using daily returns in the look-back window and rebalancing the portfolio on the last business day of each week.
Figure 5: Evolution of asset weights and asset / factor risk contributions (in %) for the Risk Budgeting strategy
In Figure 5, we observe that the asset weights are stable over time and that the risk contributions of the assets are very well aligned with the given vector of risk budgets, with the exception of small deviations which are linked to the accuracy of the Expected Shortfall estimates at a relatively high threshold $\alpha$ on a sample with a limited number of points. However, when we examine the factor risk contributions, which the RB portfolio does not take into account when optimizing portfolios, we observe that the risk contributions of Real Rates can reach levels of the order of 70%, which should preferably be avoided. On the contrary, the risk contributions from Broad Commodities are often marginal, despite the fact that 20% of risk budgets have been allocated to the commodity asset class through asset risk budgets.
To address this issue, we can consider the long-only FRB strategy that allows us to control factor risk contributions. We choose the following factor risk budgets for Growth, Real Rates, Broad Commodities and Emerging Markets respectively:
@@
The long-only FRB strategy is expected to allocate risk across the factors very closely to the selected factor risk budgets (although not perfectly in line, as weights are constrained to be positive) keeping factor risk contributions at desired levels. As shown in Figure 6, this expectation is validated, with factor risk contributions of the portfolio consistently aligning with the chosen factor risk budgets. However, in this scenario, we lose control over asset risk contributions, resulting in highly fluctuating asset weights. This outcome is undesirable for an investment strategy as it leads to high turnover and consequently increased costs.
Figure 6: Evolution of asset weights and asset / factor risk contributions (in %) for the long-only Factor Risk Budgeting strategy.
The Asset-Factor Risk Budgeting strategy can bring solutions to the issues oberserved with the two above strategies. We consider the same asset and factor risk budgets, $b_{a}$ and $b_{f}$. We also need to set the asset and factor importance parameters, which we choose equal: $\lambda_{a}=0.5$ and $\lambda_{f}=0.5$. We can now compute the AFRB portfolio by solving Problem (7) for Expected Shortfall with a threshold $\alpha=95\%$ on a weekly basis and implement the AFRB strategy.
Figure 7: Evolution of asset weights and asset / factor risk contributions (in %) for the Asset-Factor Risk Budgeting strategy.
In Figure 7, we observe that the AFRB strategy maintains a balance between asset and factor risk contributions. Compared with the case of RB, we observe a deviation from asset risk budgets. However, this is the price we have to pay to stay close to factor risk budgets. We note that the risk contribution of Broad Commodities is no longer small as it was with the RB strategy, and that we achieve a significant exposure to this factor while preserving the stability of asset weights contrary to the case of the long-only FRB. Growth and Real Rates now have balanced contributions, in line with factor risk budgets. Thanks to the importance parameters, we can choose to be more conservative in one of the two types of risk contributions, depending on the investment objective or market conditions. In short, the AFRB strategy can provide a flexible framework for better risk management of investment portfolios.
Figure 8: Cumulative performances of different strategies
Although the primary objective of these strategies is diversification rather than return maximization, it is still important to understand their behavior and performance beyond pure risk allocation figures. However, due to the limited length of our dataset, we can only analyse the strategies over a period of nearly four years which may not be sufficient to draw extensive conclusions, all the more given the major events over the period: the Covid-19 outbreak, the major increases in interest rates by central banks, periods of intense geopolitical tensions leading to war, etc.
$\hat{\mu}$ $\hat{\sigma}$ $\hat{\text{ES}}_{0.95}$ MaxDD Turnover
RB -1.42 6.55 0.93 17.12 1.00
FRB (long-only) -0.07 7.99 1.18 16.71 17.25
AFRB -0.62 7.04 1.01 16.26 2.43
Table 3: Sample mean (annualized), volatility (annualized), Expected Shortfall (daily), maximum drawdown and average turnover of the strategies (in %)
Table 3 demonstrates the main statistics associated with the different strategies computed on daily returns after transaction costs. [^22] We account for transaction costs by factoring in bid/ask spreads of 0.02% for equities, 0.01% for bonds, and 0.05% for commodities. We observe negative annualized returns from the three strategies, primarily due to their high exposure to government bonds asset class and real rates factor, which performed poorly after the end of 2021 due to the increases in interest rates. However, the long-only FRB and AFRB strategies mitigate these losses by avoiding high exposures to this factor and also by providing significant exposure to the commodity factor, a factor that performed well after the Covid-19 outbreak. In terms of management costs, the long-only FRB strategy seems inefficient, despite the fact that it is the best performer, due to its high turnover, necessitating significant changes in positions every week and resulting in high transaction costs, which, in total, amounted to 0.31% in our backtest (with 0.02% for the RB and 0.05% for the AFRB). [^23] It is important to note that the results may vary under other transaction cost schemes or rebalancing frequencies. Conversely, the AFRB strategy is a low turnover strategy similar to the RB strategy limiting the implementation cost as well as allowing for a control on exposures to the desired factors.
@@ 7 Conclusion
In this paper, we address the question of incorporating factor risk contributions into the portfolio optimization process. First, we recall the concept of Risk Budgeting, which takes into account the risk contributions of individual assets. Next, we develop the necessary tools and propose a factor risk measure that allows us to apply these ideas to underlying risk factors in the presence of a linear factor model. We then formally define Factor Risk Budgeting and an associated stochastic optimization problem for finding FRB portfolios for a large set of risk measures. Finally, we propose a framework that combines RB and FRB to obtain portfolios that control the risk induced by both assets and factors. We conclude the paper with two case studies where we construct portfolios based on our ideas using equity and macroeconomic factor models.
Acknowledgment
The authors would like to thank the participants of QuantMinds International 2023 held in London, the 68th Euro Working Group for Commodity and Financial Modelling (EWGCFM) conference held at Khalifa University in Abu Dhabi, and the XXV Workshop on Quantitative Finance held in Bologna for their valuable contributions to the discussions that occurred during these events. The quantitative teams at Amundi should also be thanked for their support. We would also like to express our gratitude to the two anonymous referees whose comments helped us to improve our article.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.